# Hard Examples Are All You Need: Maximizing GRPO Post-Training Under Annotation Budgets

### v2

@@ Abstract
Collecting high-quality training examples for language model fine-tuning is expensive, with practical budgets limiting the amount of data that can be procured. We investigate a critical question for resource-constrained alignment: under a fixed acquisition budget, should practitioners prioritize examples that are easy, medium, hard, or of random difficulty? We study Group Relative Policy Optimization (GRPO) fine-tuning across different model sizes and families, comparing four subset selection policies chosen from the same unlabeled pool using base-model difficulty estimates obtained via multi-sample evaluation. Our experiments [^2] Code available at https://github.com/Pikus16/grpo_difficulty reveal that training on the hardest examples yields the largest performance gains, up to 47%, while training on easy examples yield the smallest gains. Analysis reveals that this effect arises from harder examples providing more learnable opportunities during GRPO training. These findings provide practical guidance for budget-constrained post-training: prioritizing hard examples yields substantial performance gains on reasoning tasks when using GRPO.
@@ 1 Introduction
Large language model (LLM) alignment and post-training are fundamentally constrained by the expense of acquiring high-quality supervision data, making the selection of training examples a critical factor in achieving optimal model performance [17]. While recent advances in reinforcement learning from human feedback (RLHF) and related techniques have demonstrated remarkable improvements in model capabilities, practitioners face a practical challenge: given limited resources for data annotation and curation, which examples should be prioritized to maximize post-training performance?
@@
    Should practitioners prioritize hard examples (where the base model frequently fails), easy examples (where it frequently succeeds), middle-difficulty examples, or a random selection?
@@
Budget-aware evaluation protocol. We develop a systematic framework for comparing difficulty-conditioned training subsets under GRPO, using multi-sample base-model probing to robustly estimate example difficulty [26]. This protocol enables fair comparison across selection strategies while controlling for computational budget and other confounding factors.
Comprehensive experimental evaluation. We conduct extensive experiments on GSM8K grade-school math problems and BIG-Bench Hard’s Tracking Shuffled Objects task across models (Qwen3-4B, Qwen3-14B, Phi-4, and Llama3.1-8B) [6, 29, 1]. Our results consistently show that training on the hardest 10% of examples yields superior test performance on reasoning tasks.
Mechanistic understanding of selection effects. We analyze why hard examples prove most effective under GRPO: the algorithm requires within-group outcome variance to generate learning signals. When all samples in a group produce identical rewards, the advantages become zero and learning stops. Our analysis shows that hard examples maintain mixed outcomes longer than easy examples, which quickly converge to deterministic success [20].
Out-of-distribution generalization analysis. We evaluate models trained on different difficulty-based subsets against the AIME2025-I benchmark, a substantially harder test set than the in-distribution GSM8K [16, 2]. We find that training on the hardest examples not only improves in-distribution performance, but is also the only strategy that yields meaningful gains on OOD data.
Our findings have immediate practical implications: when using GRPO for reasoning task fine-tuning under budget constraints, practitioners should prioritize collecting and annotating examples where the base model struggles but can occasionally succeed. This hard-example focus can yield performance improvements exceeding 30 percentage points over easy selection on some configurations, transforming marginally effective fine-tuning into highly successful model improvement.
Our study isolates offline difficulty‑based selection under GRPO. It complements online difficulty‑targeted selection with rollout replay that prioritizes moderate difficulty [21], advances in process‑reward RL (PRIME) [5], and analyses of GRPO optimization biases such as Dr. GRPO [15]. We focus on outcome‑reward GRPO; extending our protocol to dense process rewards or debiased GRPO is left to future work.
@@ 2.1 Goal
Our primary research goal is to determine which subset selection strategy maximizes post-GRPO test accuracy when constrained to using only $p=10\%$ of the available training pool. Specifically, given a pool of unlabeled prompts $\mathcal{X}$, we select subsets $S\subset\mathcal{X}$ with $|S|=\lfloor p|\mathcal{X}|\rfloor$ according to different difficulty-based policies, train models using GRPO on each subset, and compare their performance on held-out test sets. This setup mirrors practical scenarios where annotation budgets limit the number of examples that can be used for fine-tuning.
@@ 2.2 Difficulty estimation via multi-sample probing
To robustly estimate example difficulty, we employ multi-sample evaluation using the base model before any fine-tuning. For each prompt $x$ in the training pool, we sample $K$ independent completions from the base model $\pi_{\text{base}}$ using temperature $\tau=1.0$ and chain-of-thought prompting where applicable [26]. For each prompt $x$, we compute:
$\hat{p}(x)$ $=\frac{1}{K}\sum_{i=1}^{K}\mathbf{1}[\text{completion}_{i}\text{ is correct}]$ (1)
@@
  • GSM8K: Exact match of the final numerical answer after extracting from the generated solution [4]
  • Tracking Shuffled Objects: Correct identification of all object positions after the described swaps [22]
@@ 2.3 Subset selection policies
    Hard(est): Select prompts with the lowest success rates:
    $S_{\text{hard}}=\arg\min_{S:|S|=\lfloor p|\mathcal{X}|\rfloor}\sum_{x\in S}\hat{p}(x)$
@@
    Middle: Select prompts nearest to the median difficulty, specifically those in the interquartile range around the median $\hat{p}(x)$ value
@@
1. Difficulty Estimation: Sample $K$ completions per prompt from base model $\rightarrow$ Compute success rate $\hat{p}(x)$
2. Subset Selection: Apply policy (Hard, Easy, Middle, Random) $\rightarrow$ Select 10% subset
3. GRPO Training: Fine-tune model on selected subset using group advantages
4. Evaluation: Test on held-out benchmark
@@ 2.4 Training procedure
We employ Group Relative Policy Optimization (GRPO) following the formulation introduced in DeepSeekMath [20]. More details on the specifics of the GRPO algorithm and hyperparameters used can be found in the appendix.
Our experiments span two benchmark datasets that probe different aspects of model capabilities. GSM8K provides 7,473 grade-school math problems requiring multi-step reasoning and arithmetic [4]. The Tracking Shuffled Objects task from BIG-Bench Hard contains 1,000 problems testing the ability to maintain state through sequences of object swaps [22]. We hold the number of prompts, rollouts per prompt, group size, KL schedule, max length, and total RL update steps constant across policies. This avoids claims being confounded by extra tokens/updates.
@@ 3.1 Hard examples drive reasoning improvements
Dataset Model Easy Middle Hard Random
@@
Figure 2: Performance by checkpoint across different subset selection policies. We see that the model trained on the hardest subset achieves superior performance by checkpoint 300 and maintains this advantage. We also see that training on easy examples performs significantly worse than all other subsets.
The results reveal striking patterns in how difficulty-based selection impacts GRPO training effectiveness. On GSM8K, the superiority of hard example selection is unambiguous across all models. The Hard policy achieves remarkable performance improvements of 34.19% for Qwen3-4B, 39.42% for Qwen3-14B, and 37.3% for Phi-4. These gains dwarf those achieved by easy selection, which manages only 3.49%, 8.26%, and 14.94% respectively. The magnitude of these differences (exceeding 30 percentage points ) suggests that training on challenging examples fundamentally alters the learning dynamics under GRPO. We see training on either middle or random difficulties yield significant gains compared to training on easy examples, but not near the gains of training on the hardest subset - affirming the hypothesis that the harder the training subset, the larger the gains we see[4].
The Tracking Shuffled Objects task amplifies these patterns. Hard selection yields positive gains of 13.60%, 22.40%, and 32.80% for Qwen3-4B, Qwen3-14B, and Phi-4 respectively. For Qwen3-4B, we see all other subsets produce either zero or negative gains. For Qwen3-14B, improvements show a stepwise pattern: training on the easy subset performs poorly, middle and random subsets provide moderate gains, and the hard subset delivers the largest gains. For Phi-4, training on the hardest subset gives the highest gains, but training on any subset still produces substantial improvements, effectively saturating the benchmark [22].
@@ 4 Analysis
We will overview the analysis on why hard examples are most effective for GRPO training on reasoning tasks in this section. We examine learning dynamics through the lens of the “learnable percentage” metric. Furthermore, we inspect performance on an out-of-distribution set to measure generalization capabilities when trained on different subsets.
@@ 4.1 Why hard examples help: maintaining learnable samples
$\text{Learnable \%}\;=\;100\cdot\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\!\left\{\,\mathsf{StdDev}(\mathbf{r}_{t})>0\,\right\},$
@@
Table 2 shows the % learnable across strategies and datasets. We see that, unsurprisingly, the hardest strategies have the highest % learnable examples, meaning training on this data subset gives the most opportunity for the model to learn during training. Conversely, we see that the easy strategy has very few learnable examples. To test whether this effect depends on the model’s baseline ability, we evaluate Llama-3.1-8B-Instruct [6], which has much lower initial performance on GSM8K. With a weaker baseline, we expect smaller performance gaps between subsets. Table 3 shows that while the hardest subset still achieves the largest gain, it is closer to the other subsets. Notably, all subsets have much higher % learnable values than the corresponding Qwen results (e.g., for the easiest subset, $21\%$ vs. $4\%$ for Qwen3-4B).
@@
easiest 9.00 3.70 2.30 7.80 14.20 6.60
middle 5.70 24.50 7.20 18.90 15.30 23.30
hardest 14.40 34.10 13.50 29.50 19.10 40.20
@@
easiest 36.54 20.90
middle 39.42 42.00
hardest 40.33 55.20
@@
Our analysis reveals why hard example selection proves dramatically more effective for GRPO training. Hard examples continue providing useful learning signal throughout training, while easy examples quickly become ”solved” and stop contributing to improvement. This sustained learning opportunity explains the remarkable performance gaps we observe (often exceeding 30 percentage points between hard and easy selection strategies).
@@ 4.2 Out-Of-Distribution Performance
Having shown that harder training subsets improve test-set performance, we further evaluate on an out-of-distribution (OOD) test set. Specifically, we take the Qwen3-4B models trained on each subset and evaluate on AIME2025-I, a dataset of 15 questions significantly more difficulty than GSM8K [2, 16]. Table 4 presents the results.
@@
Easiest 33.3%
Middle 26.7%
Hardest 40.0%
@@
We see that the training on the hardest subset is the only one to outperform the base Qwen3-4B model. Surprisingly, training on the middle subset actually does worse than training on easy or random subsets.
We see that training on the hardest subset is the only condition that meaningfully outperforms the base Qwen3-4B model, achieving a relative improvement of +20%. This reinforces the trend observed on the in-distribution test set: exposure to more challenging problems during training generalizes better to harder OOD problems. Interestingly, overall performance on the easiest and random subsets remains identical to the base model, with closer inspection showing that these subsets improved performance on certain problems but regressed performance in other areas. In contrast, training on the middle-difficulty subset yields a notable drop in performance, with no other gains.

### v3

@@ Abstract
Collecting high-quality training examples for language model fine-tuning is expensive, with practical budgets limiting the amount of data that can be procured. We investigate whether example difficulty affects GRPO training effectiveness by comparing selection strategies (easy, medium, hard, random) across multiple models and reasoning tasks. Training on the hardest 10% of examples (those where the base model fails most often) yields dramatic performance gains up to 47%, while easy examples produce minimal improvements of 3-15%. This occurs because GRPO requires outcome variance to generate learning signals; hard examples maintain mixed success/failure outcomes throughout training while easy examples quickly converge to consistent success, eliminating learning opportunities. Moreover, models trained on hard examples show superior out-of-distribution generalization, with only hard-trained models achieving meaningful gains on the AIME2025 benchmark. Our findings provide clear guidance: when budget-constrained, prioritize collecting and annotating examples where your base model struggles, as these drive nearly all learning value in GRPO fine-tuning.
@@ 1 Introduction
Large language model (LLM) alignment and post-training are fundamentally constrained by the expense of acquiring high-quality supervision data, making the selection of training examples a critical factor in achieving optimal model performance Ouyang et al. (2022). [^1] Code available here While recent advances in reinforcement learning from human feedback (RLHF) and related techniques have demonstrated remarkable improvements in model capabilities, practitioners face a practical challenge: given limited resources for data annotation and curation, which examples should be prioritized to maximize post-training performance?
@@
    Should practitioners prioritize examples that are hard (where the base model frequently fails), easy (where it frequently succeeds), medium, or a selection of random difficulty examples?
@@
  5. 5.
    Can we achieve similar performance gains by training only on examples the base model gets wrong, avoiding the circular dependency of knowing which examples will improve?
Our findings reveal a striking pattern: hard examples are all you need. Training on the hardest 10% of examples yields performance gains up to 47%, while easy examples produce minimal improvements of 3-15%. This 30+ percentage point gap transforms marginally effective fine-tuning into highly successful model improvement.
@@
Budget-aware evaluation protocol. We develop a systematic framework for comparing difficulty-conditioned training subsets under GRPO, using multi-sample base-model probing to robustly estimate example difficulty Wang et al. (2022). This protocol enables fair comparison across selection strategies while controlling for computational budget and other confounding factors.
Comprehensive experimental evaluation. We conduct extensive experiments on GSM8K grade-school math problems and BIG-Bench Hard’s Tracking Shuffled Objects task across models (Qwen3-4B, Qwen3-14B, Phi-4, and Llama3.1-8B) Dubey et al. (2024); Yang et al. (2025); Abdin et al. (2024). Our results consistently show that training on the hardest 10% of examples yields superior test performance on reasoning tasks.
Mechanistic understanding of selection effects. We analyze why hard examples prove most effective under GRPO: the algorithm requires within-group outcome variance to generate learning signals. When all samples in a group produce identical rewards, the advantages become zero and learning stops. Our analysis shows that hard examples maintain mixed outcomes longer than easy examples, which quickly converge to deterministic success Shao et al. (2024).
Out-of-distribution generalization analysis. We evaluate models trained on different difficulty-based subsets against the AIME2025-I benchmark, a substantially harder test set than the in-distribution GSM8K OpenCompass (2025); Balunović et al. (2025). We find that training on the hardest examples not only improves in-distribution performance, but is also the only strategy that yields meaningful gains on OOD data.
Base wrong vs base right analysis. We validate our “hard examples” principle from a complementary perspective by categorizing examples based on base model performance. This analysis reveals that nearly all learning value comes from examples the base model initially gets wrong, with training on “base wrong” examples achieving up to 23.5% relative improvements over “base right” examples. This convergent evidence that both low pass@k examples and base-wrong examples drive learning confirms that GRPO fundamentally benefits from training at the frontier of model capabilities.
Our findings have immediate practical implications: when using GRPO for reasoning task fine-tuning under budget constraints, practitioners should prioritize collecting and annotating examples where the base model struggles. We validate this principle from two complementary angles: examples with low pass@k success rates and examples the base model gets wrong both drive superior learning outcomes. This convergence demonstrates that “hard examples” identified through either metric capture the same underlying phenomenon: problems at the frontier of model capabilities that provide rich learning signals. This hard-example focus can yield performance improvements exceeding 30 percentage points over easy selection on some configurations, transforming marginally effective fine-tuning into highly successful model improvement.
Our study isolates offline difficulty‑based selection under GRPO. It complements online difficulty‑targeted selection with rollout replay that prioritizes moderate difficulty Sun et al. (2025), advances in process‑reward RL (PRIME) Cui et al. (2025), and analyses of GRPO optimization biases such as Dr. GRPO Liu et al. (2025). We focus on outcome‑reward GRPO; extending our protocol to dense process rewards or debiased GRPO is left to future work.
@@ 2.1 Goal
[…]
@@ 2.2 Difficulty estimation via multi-sample probing
To robustly estimate example difficulty, we employ multi-sample evaluation using the base model before any fine-tuning. For each prompt $x$ in the training pool, we sample $K$ independent completions from the base model $\pi_{\text{base}}$ using temperature $\tau=1.0$ and chain-of-thought prompting where applicable Wang et al. (2022). For each prompt $x$, we compute:
[…]
@@
  • GSM8K: Exact match of the final numerical answer after extracting from the generated solution Cobbe et al. (2021)
  • Tracking Shuffled Objects: Correct identification of all object positions after the described swaps Suzgun et al. (2023)
@@ 2.3 Subset selection policies
    Hard(-est): Select prompts with the lowest success rates:
[…]
@@
    Medium: Select prompts nearest to the median difficulty, specifically those in the interquartile range around the median $\hat{p}(x)$ value
@@
1. Difficulty Estimation: Sample $K$ completions per prompt from base model Compute success rate $\hat{p}(x)$
2. Subset Selection: Apply policy (Hard, Easy, Medium, Random) Select 10% subset 3. GRPO Training: Fine-tune model on selected subset using group advantages 4. Evaluation: Test on held-out benchmark
@@ 2.4 Training procedure
We employ Group Relative Policy Optimization (GRPO) following the formulation introduced in DeepSeekMath Shao et al. (2024). More details on the specifics of the GRPO algorithm and hyperparameters used can be found in the appendix.
Our experiments span two benchmark datasets that probe different aspects of model capabilities. GSM8K provides 7,473 grade-school math problems requiring multi-step reasoning and arithmetic Cobbe et al. (2021). The Tracking 7 Shuffled Objects task from BIG-Bench Hard contains 250 problems testing the ability to maintain state through sequences of object swaps Suzgun et al. (2023), which we randomly split 50%/50% into train/test. We hold the number of prompts, rollouts per prompt, group size, KL schedule, max length, and total RL update steps constant across policies. This avoids claims being confounded by extra tokens/updates.
@@ 3.1 Hard examples drive reasoning improvements
Dataset Model Easy Medium Hard Random
@@
Figure 2: GRPO training dynamics reveal early and persistent advantages of hard example selection. Each subplot shows test accuracy over 1000 training steps for models trained on different difficulty-based subsets (10% of full data each). The hardest subset establishes superiority by step 300 and maintains this advantage.
The results reveal striking patterns in how difficulty-based selection impacts GRPO training effectiveness. On GSM8K, the superiority of hard example selection is unambiguous across all models. The Hard policy achieves remarkable performance improvements of 34.19% for Qwen3-4B, 39.42% for Qwen3-14B, and 37.3% for Phi-4. These gains dwarf those achieved by easy selection, which manages only 3.49%, 8.26%, and 14.94% respectively. The magnitude of these differences (exceeding 30 percentage points ) suggests that training on challenging examples fundamentally alters the learning dynamics under GRPO. We see training on either medium or random difficulties yield significant gains compared to training on easy examples, but not near the gains of training on the hardest subset - affirming the hypothesis that the harder the training subset, the larger the gains we seeCobbe et al. (2021).
The Tracking Shuffled Objects task amplifies these patterns. Hard selection yields positive gains of 13.60%, 22.40%, and 32.80% for Qwen3-4B, Qwen3-14B, and Phi-4 respectively. For Qwen3-4B, we see all other subsets produce either zero or negative gains. For Qwen3-14B, improvements show a stepwise pattern: training on the easy subset performs poorly, medium and random subsets provide moderate gains, and the hard subset delivers the largest gains. For Phi-4, training on the hardest subset gives the highest gains, but training on any subset still produces substantial improvements, effectively saturating the benchmark Suzgun et al. (2023).
@@ 4 Analysis
In this section, we try to better understand why hard examples are most effective for GRPO training in three ways. First, we examine learning dynamics through the lens of new metrics - “learnable percentage”. Second, we inspect performance on an out-of-distribution set to measure generalization capabilities when trained on different subsets. Third, we look at training just on examples where the base model is wrong vs where it is right.
@@ 4.1 Why hard examples help: maintaining learnable samples
[…]
@@
Table 2 shows the % learnable across strategies and datasets. We see that, unsurprisingly, the hardest strategies have the highest % learnable examples, meaning training on this data subset gives the most opportunity for the model to learn during training. Conversely, we see that the easy strategy has very few learnable examples. To test whether this effect depends on the model’s baseline ability, we evaluate Llama-3.1-8B-Instruct Dubey et al. (2024), which has much lower initial performance on GSM8K. With a weaker baseline, we expect smaller performance gaps between subsets. Table 3 shows that while the hardest subset still achieves the largest gain, it is closer to the other subsets. Notably, all subsets have much higher % learnable values than the corresponding Qwen results (e.g., for the easiest subset, $21\%$ vs. $4\%$ for Qwen3-4B).
@@
Easy 9.00 3.70 2.30 7.80 14.20 6.60
Medium 5.70 24.50 7.20 18.90 15.30 23.30
Hard 14.40 34.10 13.50 29.50 19.10 40.20
@@
Easy 36.54 20.90
Medium 39.42 42.00
Hard 40.33 55.20
@@
Our analysis reveals why hard example selection proves dramatically more effective for GRPO training. Hard examples continue providing useful learning signal throughout training, while easy examples quickly become “solved” and stop contributing to improvement. This sustained learning opportunity explains the remarkable performance gaps we observe (often exceeding 30 percentage points between hard and easy selection strategies).
@@ 4.2 Out-Of-Distribution Performance
Having shown that harder training subsets improve test-set performance, we further evaluate on an out-of-distribution (OOD) test set. Specifically, we take the Qwen3-4B models trained on each subset and evaluate on AIME2025-I, a dataset of 15 questions significantly more difficulty than GSM8K Balunović et al. (2025); OpenCompass (2025). Table 4 presents the results.
@@
Easy 33.3%
Medium 26.7%
Hard 40.0%
@@
We see that the training on the hardest subset is the only one to outperform the base Qwen3-4B model. Surprisingly, training on the medium subset actually does worse than training on easy or random subsets.
We see that training on the hardest subset is the only condition that meaningfully outperforms the base Qwen3-4B model, achieving a relative improvement of +20%. This reinforces the trend observed on the in-distribution test set: exposure to more challenging problems during training generalizes better to harder OOD problems. Interestingly, overall performance on the easy and random subsets remains identical to the base model, with closer inspection showing that these subsets improved performance on certain problems but regressed performance in other areas. In contrast, training on the medium difficulty subset yields a notable drop in performance, with no other gains.
@@
4.3 Base Wrong vs Base Right: Decomposing GRPO Value
Having established that hard examples (those with low pass@k success rates) drive the most value in GRPO training, we now validate this principle from a complementary perspective. We investigate whether the superiority of hard examples stems from a more fundamental property: that they represent problems the base model cannot reliably solve.
We categorize training examples into four types based on base model performance (here base means the model pre GRPO training):
  1. 1.
    Base Wrong: Examples the base model gets wrong (accuracy <25% )
  2. 2.
    Base Right: Examples the base model gets right (accuracy 25% )
  3. 3.
    Base Right, Same Size As Base Wrong: A subset of the full Base Right Set that is the same size as Base Wrong, to account for model size
  4. 4.
    All: All the available examples in the train set
We hypothesize that training just on the examples where the base model is wrong will significantly match the base right subsets, as it provides the most opportunity for genuine capability improvement and higher learnable percentage.
4.3.1 Results
Table 5 presents the absolute performance after training on each subset:
ShuffleObj GSM8K
Training Set Phi-4 Qwen3-4B Qwen3-8B Phi-4 Qwen3-4B Qwen3-8B
Base Wrong 0.98 0.73 0.81 0.92 0.84 0.86
Base Right 0.98 0.64 0.83 0.80 0.70 0.72
Base Right (Same Size) 0.98 0.67 0.75 0.78 0.68 0.71
All 0.97 0.66 0.84 0.85 0.77 0.85
Table 5: Absolute performance on test sets after training on different subsets based on base model performance. Bold indicates best or tied-best performance.
Table 6 shows the relative improvements of Base Wrong training over both Base Right subsets. Table 7 shows the sizes of each subset - Base Wrong is, on average, less than half the size of Base Right.
ShuffleObj GSM8K
Metric Phi-4 Qwen3-4B Qwen3-8B Phi-4 Qwen3-4B Qwen3-8B Average
Improvement Over Base Right (Same Size)
Absolute 0.0% 6.0% 5.8% 14.0% 16.0% 15.0% 10.2%
Relative 0.0% 9.0% 7.7% 17.9% 23.5% 21.1% 14.3%
Improvement Over Base Right (Full)
Absolute 0.0% 9.0% -2.0% 12.0% 14.0% 14.0% 7.8%
Relative 0.0% 14.1% -2.4% 15.0% 20.0% 19.4% 11.0%
Table 6: Improvements from training on Base Wrong examples compared to Base Right examples. Relative improvement calculated as (Base Wrong - Base Right) / Base Right.
4.3.2 Analysis
The results strongly support our hypothesis that GRPO learning value concentrates in examples the base model gets wrong:
Base Wrong consistently outperforms Base Right: Across models and tasks, training exclusively on Base Wrong examples achieves equal or superior performance to training on Base Right, even when Base Right has significantly more examples. The average relative improvement of 14.3% over size-matched Base Right subsets demonstrates that these examples provide fundamentally different learning signals. Even when we train on the full Base Right set, we see that Base Wrong outperforms it by 11% on average.
Base Wrong matches / outperforms All: Training on all Base Wrong samples actually slightly outperforms training on All samples on average (85% vs 82%), further supporting that most, if not all, meaningful training signal is concentrated in the Base Wrong subset.
Task-specific patterns emerge: On GSM8K, the advantage is particularly pronounced, with Base Wrong training achieving 15-23.5% relative improvements. This aligns with our earlier findings that mathematical reasoning benefits most from challenging examples. ShuffleObj shows more modest gains, potentially due to the discrete nature of object tracking limiting the diversity of solution paths.
Model-specific sensitivity: Phi-4 shows less differentiation between strategies on ShuffleObj (all achieving 0.98), suggesting potential ceiling effects. However, on GSM8K, even Phi-4 shows substantial 17.9% relative improvement, confirming the generality of the phenomenon.
Practical implications: These findings provide a practical recipe for GRPO dataset curation: practitioners can identify valuable training examples by simply evaluating base model performance, without needing to know which examples will ultimately improve. This avoids the circular dependency inherent in selecting “learnable” examples and provides a straightforward criterion for data collection.
ShuffleObj GSM8K
Training Set Phi-4 Qwen3-4B Qwen3-8B Phi-4 Qwen3-4B Qwen3-8B
Base Wrong 17 37 13 343 409 371
Base Right 108 88 112 657 591 629
All 125 125 125 1000 1000 1000
Table 7: Subset sizes for each dataset, model, and training strategy. On average, Base Wrong is only 42% the size of Base Right, and yet consistently significantly outperforms this subset. Note that GSM8K is clipped to 1000 since we only run for 1000 steps with 1 prompt per update.
The concentration of learning value in Base Wrong examples validates our central thesis from a complementary angle. Our two experimental approaches converge on the same insight: examples with low pass@k rates (our “hardest” selection) are predominantly those where the base model fails (“base wrong”), while examples with high pass@k rates (our “easy” selection) are those the base model already masters (“base right”). This convergence demonstrates that whether we measure difficulty through multi-sample success rates or binary base model performance, we identify the same high-value training examples: those at the frontier of model capabilities where learning can actually occur.